An aeroplane heads due north at 500 km/h. It experiences a 80 km/h crosswind flowing in


the direction N60^°E.


(a) Find the true velocity of the aeroplane.

1
Expert's answer
2022-04-29T14:03:03-0400
"\vec{v}=500\vec{j}"

"\vec{u}=80\sin60\degree \vec{i}+80\cos60\degree\vec{j}"

"\vec{v}_{res}=\vec{v}+\vec{u}"

"=40\sqrt{3}\vec{i}+540\vec{j}"

"|\vec{v}_{res}|=\sqrt{(40\sqrt{3})^2+(540)^2}=20\sqrt{741}(km/h)"

"\tan \theta=\dfrac{40\sqrt{3}}{540}=\dfrac{2\sqrt{3}}{27}"

"\theta=\tan^{-1}\dfrac{2\sqrt{3}}{27}\approx7.3111\degree"

"20\sqrt{741}" km/h in the direction "N(\tan^{-1}\dfrac{2\sqrt{3}}{27})\degree E."


"544.4263" km/h in the direction "N(7.31)\degree E."